Jack holds a portfolio with the following securities: Security Investment Return Stock A 639,741 -1.3% Stock B 378,915 4.0% Stock C 741,180 9.0% Calculate the expected return of portfolio. Round the answers to two decimal places in percentage form.
Answer #1

Total investment=(639741+378915+741180)=$1759836

Expected return of portfolio=Respective return*Respective investment weight

=(639741/1759836*-1.3)+(378915/1759836*4)+(741180/1759836*9)

which is equal to

=4.18%(Approx).